# Inflations among quantum Grothendieck rings of type A

### v2

@@ 1. Introduction
This result partly motivates us to investigate further relationship among the quantum Grothendieck rings. In this paper, we restrict our attention to type $\mathrm{A}$ and construct a collection of injective homomorphisms among the quantum Grothendieck rings of different ranks. (Of course, by composing them with the isomorphisms between type $\mathrm{B}_{n}$ and type $\mathrm{A}_{2n-1}$ in [6, 7], we can extend our collection to a collection of injective homomorphisms among type $\mathrm{AB}$.) To be more precise, let $\mathscr{C}_{n}$ be a monoidal skeleton of the category $\mathscr{C}$ (= the subcategory $\mathscr{C}_{\mathbb{Z}}$ in the sense of Hernandez–Leclerc [9]) for the quantum loop algebra of $\mathfrak{sl}_{n}$, and $K_{t}(\mathscr{C}_{n})$ its quantum Grothendieck ring. Let $[1,n]\coloneqq\{1,2,\ldots,n\}$ be the integer interval between $1$ and $n$, and $I_{n}\coloneqq[1,n-1]$ an index set of simple roots of $\mathfrak{sl}_{n}$. The main result of this paper is the following.
@@
Let $n,{\tilde{n}}$ be two positive integers with $1<n<{\tilde{n}}$. For any choice of height functions $\xi\colon I_{n}\to\mathbb{Z}$, ${\tilde{\xi}}\colon I_{\tilde{n}}\to\mathbb{Z}$ (cf. §3.3) and a (strictly) increasing function $\nu\colon[1,n]\to[1,{\tilde{n}}]$, we have an injective homomorphism $\Psi_{{\tilde{\xi}},\nu,\xi}\colon K_{t}(\mathscr{C}_{n})\to K_{t}(\mathscr{C}_{\tilde{n}})$ respecting the canonical bases.
We call the homomorphisms $\Psi_{{\tilde{\xi}},\nu,\xi}$ the quantum inflations, as the classical limit of $\Psi_{{\tilde{\xi}},\nu,\xi}$ for a special choice of $(\xi,{\tilde{\xi}},\nu)$ coincides with the inflation from $K(\mathscr{C}_{n})$ to $K(\mathscr{C}_{\tilde{n}})$ studied by Brito–Chari [1, §3]. Our construction of $\Psi_{{\tilde{\xi}},\nu,\xi}$ relies on the presentation of a localization of $K_{t}(\mathscr{C}_{n})$ (resp. $K_{t}(\mathscr{C}_{\tilde{n}})$) due to Hernandez–Leclerc [11] as the bosonic extension of the half of the quantized enveloping algebra $U_{t}(\mathfrak{sl}_{n})$ (resp. $U_{t}(\mathfrak{sl}_{\tilde{n}})$), which depends on the choice of height function $\xi$ (resp. ${\tilde{\xi}}$). The choice of increasing function $\nu$ gives rise to an injective homomorphism between the bosonic extensions, which is an analog of the homomorphism $\mathfrak{sl}_{n}\to\mathfrak{sl}_{\tilde{n}}$ of Lie algebras sending the $(i,j)$-matrix unit $E_{i,j}\in\mathfrak{sl}_{n}$, $i\neq j$, to $E_{\nu(i),\nu(j)}\in\mathfrak{sl}_{\tilde{n}}$.
@@ 2.1. Notation
Throughout the paper, we use the following conventions. For a mathematical statement $P$, we set $\delta(P)$ to be $1$ or 00 according that $P$ is true or false. As a special case, we set $\delta_{i,j}\coloneqq\delta(i=j)$. We denote by $\mathbb{Z},\mathbb{Q}$, and $\mathbb{C}$ the sets of integers, rational numbers, and complex numbers respectively. For $a,b\in\mathbb{Z}$ with $a\leq b$, we set $[a,b]\coloneqq\{k\in\mathbb{Z}\mid a\leq k\leq b\}$. Given $n\in\mathbb{Z}$, we write $a\equiv_{n}b$ when $a-b\in n\mathbb{Z}$.
@@ 2.2. Quantum loop algebras and $q$-characters
Then we consider the category $\mathscr{C}_{n}$ of finite-dimensional $U_{q}(L\mathfrak{sl}_{n})$-modules whose composition factors are isomorphic to $L(m)$’s for $m\in\mathcal{M}_{n}^{+}$. This category is the same as Hernandez–Leclerc’s category $\mathscr{C}_{\mathbb{Z}}$ for $U_{q}(L\mathfrak{sl}_{n})$ in [9]. The category $\mathscr{C}_{n}$ is closed under taking tensor products and duals. Therefore, it carries a natural structure of rigid monoidal abelian category.
@@ 3.1. Definition
By construction, $K_{t}(\mathscr{C}_{n})$ is a $\mathbb{Z}[t^{\pm 1/2}]$-subalgebra of $\mathcal{Y}_{n,t}$. By (2.1), we have
@@ 3.2. Canonical bases
In [8], the standard basis element $E_{t}(m)$ is defined in terms of the element $F_{t}(Y_{i,p})$ such that $Y_{i,p}$ is the unique dominant monomial occurring in $F_{t}(Y_{i,p})$, instead of $\chi_{q}(L(Y_{i,p}))$. Our definition of $E_{t}(m)$ is equivalent to the one in loc. cit. as $F_{t}(Y_{i,p})=\chi_{q}(L(Y_{i,p}))$ holds in type $\mathrm{A}$.
@@ 3.3. Presentations
    $\phi_{\xi}(i,p\pm 2)=\begin{cases}(\tau_{\xi}^{\pm 1}(\alpha),k)&\text{if $\tau_{\xi}^{\pm 1}(\alpha)\in R^{+}_{n}$},\\ (-\tau_{\xi}^{\pm 1}(\alpha),k\pm 1)&\text{if $\tau_{\xi}^{\pm 1}(\alpha)\not\in R^{+}_{n}$}.\end{cases}$
@@
$\mathscr{N}(Y_{i,p},Y_{j,s})=(-1)^{k+l+\delta(k\geq l)}(\alpha,\beta),$
where $(\alpha,k)=\phi_{\xi}(i,p)$ and $(\beta,l)=\phi_{\xi}(j,s)$.
@@ 4. Quantum inflations
In this section, we introduce a collection of injective homomorphisms from $K_{t}(\mathscr{C}_{n})$ to $K_{t}(\mathscr{C}_{\tilde{n}})$ with $n<{\tilde{n}}$, which we call the quantum inflations as it includes a quantum analog of Brito–Chari’s inflation in §2.3 as a special case. Our main theorem (= Theorem 4.6) in §4.2 asserts that they respect the canonical bases, and hence their classical limits respect the simple classes. In particular, we obtain in §4.3 a proof of Conjecture 2.1.
@@ 4.1. Construction
We have $e(\alpha_{i})=e_{i}$ for any $i\in I_{n}$ by definition. Note that $e(\alpha_{a,b})$ and $e(\alpha_{c,d})$ commute if $b+1<c$.
@@
For $x,y,z\in U_{n}^{+}$, we have $[[x,y]_{t^{1/2}},z]_{t^{1/2}}=[x,[y,z]_{t^{1/2}}]_{t^{1/2}}$ if $xz=zx$. With this remark in mind, the assertion follows by induction on $c-a$. ∎
@@
$\sigma_{l}(e_{i,k})$ =$=\begin{cases}(t-t^{-1})^{-1}[e_{i,k},e_{l,k}]_{t^{1/2}}&\text{if$|i-l|=1$},\\ e_{i,k+\delta_{i,l}}&\text{otherwise},\end{cases}$
$\sigma^{-1}_{l}(e_{i,k})$ =$=\begin{cases}(t-t^{-1})^{-1}[e_{l,k},e_{i,k}]_{t^{1/2}}&\text{if$|i-l|=1$},\\ e_{i,k-\delta_{i,l}}&\text{otherwise}.\end{cases}$
@@ 4.2. Compatibility with canonical bases
The quantum inflation $\Psi_{{\tilde{\xi}},\nu,\xi}$ in Definition 4.3 induces an injective map from the canonical basis of $K_{t}(\mathscr{C}_{n})$ into the canonical basis of $K_{t}(\mathscr{C}_{\tilde{n}})$. In other words, there is an injective map $\psi_{{\tilde{\xi}},\nu,\xi}\colon\mathcal{M}_{n}^{+}\to\mathcal{M}_{\tilde{n}}^{+}$ such that we have
@@
$\Psi_{{\tilde{\xi}},\nu,\xi}\colon K_{t}(\mathscr{C}_{n})\hookrightarrow K_{t}(\mathscr{C}_{\tilde{n}}).$
@@
Recall that a height function $\xi\colon I_{n}\to\mathbb{Z}$ defines a Dynkin quiver $Q_{\xi}$ of type $\mathrm{A}_{n-1}$ so that we have an arrow $i\to j$ in $Q_{\xi}$ if and only if $|i-j|=1$ and $\xi(i)>\xi(j)$. For example, when $\xi$ is increasing, $Q_{\xi}$ is a monotone quiver: $Q_{\xi}=(1\leftarrow 2\leftarrow\cdots\leftarrow n-1)$. Associated to a Dynkin quiver $Q$ of type $\mathrm{A}_{n-1}$, the rescaled dual PBW generator $\tilde{E}^{*}_{Q}(\alpha)\in U_{n}^{+}$ is defined for each positive root $\alpha\in R_{n}^{+}$ as in [11, §6.1]. (Note that $\tilde{E}^{*}_{Q}(\alpha)$ is written simply as $\tilde{E}^{*}(\alpha)$ in loc. cit. and our $t$ is equal to $v$ in loc. cit.) For an arbitrary height function $\xi$, by the construction of the isomorphism $\Psi_{\xi}$ and [11, Theorem 6.1], we know that the isomorphism $\Psi_{\xi}\colon\widehat{\mathcal{A}}_{n}\to K_{t}(\mathscr{C}_{n})_{loc}$ sends $\iota_{k}(\tilde{E}^{*}_{Q_{\xi}}(\alpha))$ to $\chi_{q,t}(L(Y_{i,p}))$ with $(i,p)=\phi_{\xi}^{-1}(\alpha,k)$ for any $(\alpha,k)\in R_{n}^{+}\times\mathbb{Z}$.
@@
By Theorem 4.6, it makes sense to consider the specialization $\Psi_{{\tilde{\xi}},\nu,\xi}|_{t=1}$ at $t^{1/2}=1$ of the homomorphism $\Psi_{{\tilde{\xi}},\nu,\xi}$. It is a unique ring homomorphism from $K(\mathscr{C}_{n})$ to $K(\mathscr{C}_{\tilde{n}})$ making the following diagram commute:
@@
Recall Brito–Chari’s inflation $\Psi_{{\tilde{n}},n}\colon K(\mathscr{C}_{n})\to K(\mathscr{C}_{\tilde{n}})$ and the monoid homomorphism $\psi_{{\tilde{n}},n}\colon\mathcal{M}_{n}^{+}\to\mathcal{M}_{\tilde{n}}^{+}$ from §2.3. The next proposition, together with Corollary 4.9, verifies Conjecture 2.1.
@@ 5.1. Quiver Hecke algebras of type $\mathrm{A}_{\infty}$
For each positive root $\alpha\in R^{+}$, there is a self-dual one-dimensional module $L(\alpha)\in\mathscr{A}(\alpha)$. In [13], $L(\alpha_{a,b})$ was denoted by $L(a,b-1)$. The set $\{[L(\alpha)]\mid\alpha\in R^{+}\}$ generates the $\mathbb{Z}[t^{\pm 1/2}]$-algebra $K(\mathscr{A})$ and satisfies the relation
@@ 5.2. Categorification of $K_{t}(\mathscr{C}_{n})$
As our base field $\mathbb{C}$ is of characteristic zero, the basis of the self-dual simple modules of $K(\mathscr{A})$ correspond to the dual canonical basis of the quantum unipotent coordinate ring of type $\mathrm{A}_{\infty}$ by [21, 23]. In particular, the self-dual simple classes are parametrized by the set $(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$ (equivalent to the set of Zelevinsky’s multi-segments) as follows. Define a total ordering $\leq$ of the set $R^{+}$ so that we have $\alpha_{a,b}<\alpha_{a^{\prime},b^{\prime}}$ if $a<a^{\prime}$, or $a=a^{\prime}$ $b<b^{\prime}$. For each $\mathbf{c}=(c_{\alpha})_{\alpha\in R^{+}}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$, we define
@@
where $\overset{\leftarrow}{\circ}$ means the ordered convolution product along the above ordering of $R^{+}$ increasing from the right to left. Then, the self-dual simple class $[L(\mathbf{c})]$ in $K(\mathscr{A})$ corresponding to $\mathbf{c}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$ is characterized by the property
(5.7) $[L(\mathbf{c})]-[M(\mathbf{c})]\in\sum_{\mathrm{wt}(\mathbf{c}^{\prime})=\mathrm{wt}(\mathbf{c})}t^{-1}\mathbb{Z}[t^{-1}][M(\mathbf{c}^{\prime})],$
@@
Let $\sigma$ be the involution of $K(\mathscr{A}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]$ given by $\sigma(t^{1/2})=t^{-1/2}$ and $\sigma([M])=t^{(r_{n}(\beta),r_{n}(\beta))/2}[M^{*}]$ for $M\in\mathscr{A}_{\beta}$. The equation (5.4) tells us that $\sigma$ is an anti-involution of $K(\mathscr{A}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]$. For $M\in\mathscr{A}_{\beta}$, we define its rescaled class $[\widetilde{M}]\coloneqq t^{(r_{n}(\beta),r_{n}(\beta))/4}[M]$ so that $[\widetilde{M}]$ is fixed by $\sigma$ if and only if $[M]=[M^{*}]$. We have
$[\widetilde{L(\alpha_{a,b})}]=t^{(1-\delta_{b-a,n})/2}[L(\alpha_{a,b})]$
for any $a,b\in\mathbb{Z}$ with $a<b$. Letting
@@
we have ()¯$\overline{(\cdot)}\circ\Phi^{\prime}_{n}=\Phi^{\prime}_{n}\circ\sigma$ by Theorem 5.1. In particular, $\Phi^{\prime}_{n}([\widetilde{L(\mathbf{c})}])$ is bar-invariant for any $\mathbf{c}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$.
@@
$[\widetilde{M(\mathbf{c})}]=t^{\sum_{\alpha>\alpha^{\prime}}c_{\alpha}c_{\alpha^{\prime}}(r_{n}(\alpha),r_{n}(\alpha^{\prime}))/2}\mathop{\star}^{\leftarrow}_{\alpha\in R^{+}}[\widetilde{L(\alpha)}]^{\star c_{\alpha}}$
for any $\mathbf{c}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$ such that $c_{\alpha_{a,b}}=0$ if $b-a>n$. We also note that
@@
for any $\mathbf{c}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$, where $m_{\mathbf{c}}\coloneqq\prod_{a,b\in\mathbb{Z},b-a\in I_{n}}Y_{b-a,a+b}^{c_{\alpha_{a,b}}}\in\mathcal{M}_{n}^{+}$ (recall the definition of $E_{t}(m)$ from (3.1)). Together with the equality (5.7) and the bar-invariance of $\Phi^{\prime}_{n}([\widetilde{L(\mathbf{c})}])$, we find that, if $\Phi^{\prime}_{n}([\widetilde{L(\mathbf{c})}])$ is non-zero, it satisfies the two characterizing properties of $\chi_{q,t}(L(m_{\mathbf{c}}))$ in Theorem 3.3, and hence $\Phi^{\prime}_{n}([\widetilde{L(\mathbf{c})}])=\chi_{q,t}(L(m_{\mathbf{c}}))$. Since every simple object of $\mathscr{T}_{n}$ is the image under $\Omega_{n}$ of a simple module in $\mathscr{A}_{n}$, it verifies the assertion. ∎

### v3

@@ 1. Introduction
This result partly motivates us to investigate further relationship among the quantum Grothendieck rings. In this paper, we restrict our attention to type $\mathrm{A}$ and construct a collection of injective homomorphisms among the quantum Grothendieck rings of different ranks. (Of course, by composing them with the isomorphisms between type $\mathrm{B}_{n}$ and type $\mathrm{A}_{2n-1}$ in [6, 7], we can extend our collection to a collection of injective homomorphisms among type $\mathrm{AB}$.) To be more precise, let $\mathscr{C}_{n}$ be a monoidal skeleton of the category $\mathscr{C}$ (= the subcategory $\mathscr{C}_{\mathbb{Z}}$ in the sense of Hernandez–Leclerc [9]) for the quantum loop algebra of $\mathfrak{sl}_{n}$, and $K_{t}(\mathscr{C}_{n})$ its quantum Grothendieck ring. Let $1,n]\coloneqq\{1,2,\ldots,n\}$ be the integer interval between $1$ and $n$, and $I_{n}\coloneqq[1,n-1]$ an index set of simple roots of $\mathfrak{sl}_{n}$. The main result of this paper is the following.
@@
[…]
@@ 2.1. Notation
Throughout the paper, we use the following conventions. For a mathematical statement $P$, we set $\delta(P)$ to be $1$ or 0 according that $P$ is true or false. As a special case, we set $\delta_{i,j}\coloneqq\delta(i=j)$. We denote by $\mathbb{Z},\mathbb{Q}$, and $\mathbb{C}$ the sets of integers, rational numbers, and complex numbers respectively. For $a,b\in\mathbb{Z}$ with $a\leq b$, we set $a,b]\coloneqq\{k\in\mathbb{Z}\mid a\leq k\leq b\}$. Given $n\in\mathbb{Z}$, we write $a\equiv_{n}b$ when $a-b\in n\mathbb{Z}$.
@@ 2.2. Quantum loop algebras and $q$-characters
Then we consider the category $\mathscr{C}_{n}$ of finite-dimensional $U_{q}(L\mathfrak{sl}_{n})$-modules whose composition factors are isomorphic to $L(m)$’s for $m\in\mathcal{M}_{n}^{+}$. This category is the same as Hernandez–Leclerc’s category $\mathscr{C}_{\mathbb{Z}}$ for $U_{q}(L\mathfrak{sl}_{n})$ in [9]. (Note that the symbol $\mathscr{C}_{n}$ has a different meaning in [9].) The category $\mathscr{C}_{n}$ is closed under taking tensor products and duals. Therefore, it carries a natural structure of rigid monoidal abelian category.
@@ 3.1. Definition
Compare to (2.1). By construction, $K_{t}(\mathscr{C}_{n})$ is a $\mathbb{Z}[t^{\pm 1/2}]$-subalgebra of $\mathcal{Y}_{n,t}$. By [20, Corollary 3.6] or [8, Theorem 6.2], we have
@@ 3.2. Canonical bases
In [8], the standard basis element $E_{t}(m)$ is defined in terms of the element $F_{t}(Y_{i,p})$ such that $Y_{i,p}$ is the unique dominant monomial occurring in $F_{t}(Y_{i,p})$, instead of $\chi_{q}(L(Y_{i,p}))$. Our definition of $E_{t}(m)$ is equivalent to the one in loc. cit. as $F_{t}(Y_{i,p})=\chi_{q}(L(Y_{i,p}))$ holds in type $\mathrm{A}$. Note that the last equality is not true in general.
@@
Remark 3.4.
The transition matrix from the standard basis to the canonical basis turns out to be unitriangular with respect to the so-called Nakajima partial ordering of dominant monomials. In fact, the original defining condition of the canonical basis in [20, 8] requires this unitriangularity as well. The weaker condition in Theorem 3.3 is enough to characterize the basis, as explained in [12, Remark 7.8].
@@ 3.3. Presentations
[…]
@@
$\mathscr{N}(Y_{i,p},Y_{j,s})=(-1)^{k+l+\delta(p\geq s)}(\alpha,\beta),$
where $(\alpha,k)=\phi_{\xi}(i,p)$ and $(\beta,l)=\phi_{\xi}(j,s)$. Moreover, we have
$\mathscr{N}(Y_{i,p},Y_{j,s})=0\quad\text{if $p\geq s$ and $k<l$.}$
@@ 4. Quantum inflations
[…]
@@ 4.1. Construction
We have $e(\alpha_{i})=e_{i}$ for any $i\in I_{n}$ by definition. Note that $e(\alpha_{a,b})$ and $e(\alpha_{c,d})$ commute if $b<c$.
@@
For $x,y,z\in U_{n}^{+}$, we have $x,y]_{t^{1/2}},z]_{t^{1/2}}=[x,[y,z]_{t^{1/2}}]_{t^{1/2}}$ if $xz=zx$. With this remark in mind, the assertion follows by induction on cac-a. ∎
@@
Example 4.2.
For example, when $n=2$ and ${\tilde{n}}=3$, there are three increasing functions $\nu\colon[1,2]\to[1,3]$. The assignment (4.2) is written as
$e_{1,k}\mapsto\begin{cases}e_{1,k}&\text{if $(\nu(1),\nu(2))=(1,2)$},\\ (t-t^{-1})^{-1}[e_{1,k},e_{2,k}]_{t^{1/2}}&\text{if $(\nu(1),\nu(2))=(1,3)$},\\ e_{2,k}&\text{if $(\nu(1),\nu(2))=(2,3)$},\end{cases}$
for each $k\in\mathbb{Z}$.
@@
$\sigma_{l}(e_{i,k})$ $=\begin{cases}(t-t^{-1})^{-1}[e_{i,k},e_{l,k}]_{t^{1/2}}&\text{if $|i-l|=1$},\\ e_{i,k+\delta_{i,l}}&\text{otherwise},\end{cases}$
$\sigma^{-1}_{l}(e_{i,k})$ $=\begin{cases}(t-t^{-1})^{-1}[e_{l,k},e_{i,k}]_{t^{1/2}}&\text{if $|i-l|=1$},\\ e_{i,k-\delta_{i,l}}&\text{otherwise}.\end{cases}$
@@ 4.2. Compatibility with canonical bases
[…]
@@
[…]
compatible with the bar-involutions.
@@
Recall that a height function $\xi\colon I_{n}\to\mathbb{Z}$ defines a Dynkin quiver $Q_{\xi}$ of type $\mathrm{A}_{n-1}$ so that we have an arrow $i\to j$ in $Q_{\xi}$ if and only if |ij|=1|i-j|=1 and $\xi(i)>\xi(j)$. For example, when $\xi$ is increasing, $Q_{\xi}$ is a monotone quiver: $Q_{\xi}=(1\leftarrow 2\leftarrow\cdots\leftarrow n-1)$. Associated to a Dynkin quiver $Q$ of type $\mathrm{A}_{n-1}$, the rescaled dual PBW generator $\tilde{E}^{*}_{Q}(\alpha)\in U_{n}^{+}$ is defined for each positive root $\alpha\in R_{n}^{+}$ as in [11, §6.1]. (Note that $\tilde{E}^{*}_{Q}(\alpha)$ is written simply as $\tilde{E}^{*}(\alpha)$ in loc. cit. and our $t$ is equal to $v$ in loc. cit.) For an arbitrary height function $\xi$, by the construction of the isomorphism $\Psi_{\xi}$ and [11, Theorem 6.1], we know that the isomorphism $\Psi_{\xi}\colon\widehat{\mathcal{A}}_{n}\to K_{t}(\mathscr{C}_{n})_{loc}$ sends $\iota_{k}(\tilde{E}^{*}_{Q_{\xi}}(\alpha))$ to $\chi_{q,t}(L(Y_{i,p}))$ with $(i,p)=\phi_{\xi}^{-1}(\alpha,k)$ for any $(\alpha,k)\in R_{n}^{+}\times\mathbb{Z}$.
@@
(4.7) ${K_{t}(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}},\nu,\xi}}$$\scriptstyle{\mathop{\mathrm{ev}}\nolimits_{t=1}}$${K_{t}(\mathscr{C}_{\tilde{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathop{\mathrm{ev}}\nolimits_{t=1}}$${K(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}},\nu,\xi}|_{t=1}}$${K(\mathscr{C}_{\tilde{n}}).}$
@@
[…]
@@ 5.1. Quiver Hecke algebras of type $\mathrm{A}_{\infty}$
For each positive root $\alpha\in R^{+}$, there is a self-dual one-dimensional module $L(\alpha)\in\mathscr{A}(\alpha)$. In [13], $L(\alpha_{a,b})$ was denoted by L(a,b1)L(a,b-1). The set $\{[L(\alpha)]\mid\alpha\in R^{+}\}$ generates the $\mathbb{Z}[t^{\pm 1/2}]$-algebra $K(\mathscr{A})$ and satisfies the relation
@@ 5.2. Categorification of $K_{t}(\mathscr{C}_{n})$
As our base field $\mathbb{C}$ is of characteristic zero, the basis of the self-dual simple modules of $K(\mathscr{A})$ correspond to the dual canonical basis of the quantum unipotent coordinate ring of type $\mathrm{A}_{\infty}$ by [21, 23]. In particular, the self-dual simple classes are parametrized by the set $(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$ (equivalent to the set of Zelevinsky’s multi-segments) as follows. Define a total ordering $\leq$ of the set $R^{+}$ so that we have $\alpha_{a,b}<\alpha_{a^{\prime},b^{\prime}}$ if $a<a^{\prime}$, or $a=a^{\prime}$ & $b<b^{\prime}$. For each $\mathbf{c}=(c_{\alpha})_{\alpha\in R^{+}}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$, we define
@@
[…]
(5.7) $L(\mathbf{c})]-[M(\mathbf{c})]\in\sum_{\mathrm{wt}(\mathbf{c}^{\prime})=\mathrm{wt}(\mathbf{c})}t^{-1}\mathbb{Z}[t^{-1}][M(\mathbf{c}^{\prime})],$
@@
Let $\sigma$ be the involution of $K(\mathscr{A}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]$ given by $\sigma(t^{1/2})=t^{-1/2}$ and $\sigma([M])=t^{(r_{n}(\beta),r_{n}(\beta))/2}[M^{*}]$ for $M\in\mathscr{A}_{\beta}$. The equation (5.4) tells us that $\sigma$ is an anti-involution of $K(\mathscr{A}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]$. For $M\in\mathscr{A}_{\beta}$, we define its rescaled class $\widetilde{M}]\coloneqq t^{(r_{n}(\beta),r_{n}(\beta))/4}[M]$ so that $\widetilde{M}$] is fixed by $\sigma$ if and only if $M]=[M^{*}]$. We have
$\widetilde{L(\alpha_{a,b})}]=t^{(1-\delta_{b-a,n})/2}[L(\alpha_{a,b})]$
for any $a,b\in\mathbb{Z}$ with $b-a\in[1,n]$. Letting
@@
[…]
@@
$\widetilde{M(\mathbf{c})}]=t^{\sum_{\alpha>\alpha^{\prime}}c_{\alpha}c_{\alpha^{\prime}}(r_{n}(\alpha),r_{n}(\alpha^{\prime}))/2}\mathop{\star}^{\leftarrow}_{\alpha\in R^{+}}[\widetilde{L(\alpha)}]^{\star c_{\alpha}}$
for any $\mathbf{c}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$ such that $c_{\alpha_{a,b}}=0$ if ba>nb-a>n. We also note that
@@
[…]
Remark 5.3.
As far as the author understands, such a categorification of the quantum Grothendieck ring is known only for type $\mathrm{A}$ at this moment. One of the special features of type $\mathrm{A}$ is that the bozonic extension algebra $\widehat{\mathcal{A}}_{n}$ can be obtained as an explicit quotient of the half of the quantized enveloping algebra of type $\mathrm{A}_{\infty}$, which enables us to construct the categorification out of the quiver Hecke algebras of type $\mathrm{A}_{\infty}$. It would be very interesting if we have a similar categorification for the other types.